OpenStudy (anonymous):

How do I write this in slope intercept form? Y=3/2(x-4)+5

OpenStudy (anonymous):

distribute \[y = \frac{3}{2}x - (\frac{3}{2})(4) + 5\] \[y = \frac{3}{2} x - 6 + 5\] \[y = \frac{3}{2} x - 1\]

OpenStudy (anonymous):

Thanks